OpenStudy (anonymous):

Find f'(x) f(x)=x^2+7/3x-1

OpenStudy (campbell_st):

needs the quotient rule u = x^2 + 7 v = 3x -1 du/dx = 2x dv/dx = 3 Quotient rule is \[dy/dx = (v \times u' - u \times v')/v^2\] I'll let you finish it

OpenStudy (anonymous):

I got the answer 3x^2-2x-21/(3x-1)^2 Is this correct?

OpenStudy (anonymous):

yes.